OpenStudy (anonymous):

Please help!(: Deals with the SSA case, what would be all the possible solutions? maybe include a rough triangle sketch if possible for visual? A=50 degrees, a=10, b=15.

OpenStudy (anonymous):



OpenStudy (anonymous):

that is my rough sketch, although it is not clear that \(10\) is long enough we can solve for the angle \(B\) via \[\frac{\sin(B)}{15}=\frac{\sin(50)}{10}\] or \[\sin(B)=\frac{15\sin(50)}{10}\] or even \[\sin(B)=\frac{3\sin(50)}{2}\]

OpenStudy (anonymous):

unfortunately \(\frac{3\sin(50)}{2}>1\) so there is not such sine, and hence to such \(B\) or triangle picture sort of hinted at it

OpenStudy (anonymous):

Okay, thanks for the help(: Always appreciate it!